A computer is marked 230 V; 950 W

Calculate the resistance of the computer.


1
Expert's answer
2021-09-09T10:50:05-0400

We have to use Ohm's law to find the computer resistance:

"V=IR; P=IV \implies P = \cfrac{V^2}{R} 
\\ \implies R=\cfrac{V^2}{P}=\cfrac{(230\,V)^2}{950\,W}=55.68\,\Omega"


In conclusion, the resistance of the computer is "55.68\,\Omega".


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.